Good Morning
I was wondering if you could help me with this problem
Given F(x)=x^2+2 , Find F(a+2)-F(a)
I did
(a+2) (a+2)
a^2+4a+4
a^2+4a+4-a^2
Answer 4a+4
I'm not sure if this is correct though.
Thank you

Answer by stanbon(75887)

was wondering if you could help me with this problem
Given F(x)=x^2+2 , Find F(a+2)-F(a)
----------------------------------------
F(a+2) = (a+2)^2+2
F(a+2) = a^2+4a+4+2
F(a+2) = a^2+4a+6
---------
F(a) = a^2+2
---------
Therefore:
F(a+2)-F(a) = [a^2+4a+6]-[a^2+2] = 4a+4
==========================================
Cheers,
Stan H.